OpenStudy (anonymous):

write the equation of the line that passes through (1,3) and (4,4) in standard form.

OpenStudy (anonymous):

Okay so do you know what the standard form looks like?

OpenStudy (anonymous):

Okay well I'm leaving but I will write down some formulas that will allow you to find the equation

OpenStudy (anonymous):

Standard form is Ax^2 + Bx + C To find slope it is m = (y2 - y1)/(x2 -x1) After you find that input one of the points into the equation : y=mx+b then convert to standard form